OpenStudy (anonymous):

I need 10th grade math help

OpenStudy (anonymous):

okay...? Anything in particular?

OpenStudy (anonymous):

ABCD is a rectangle. Find the length of each diagonal. AC=2(x-3) BD=x+5

Vocaloid (vocaloid):

AC and BD are the diagonals, and diagonals of a rectangle are equal to each other 2(x-3) = x + 5 the first step is to solve for x, can you do that for me please?

OpenStudy (anonymous):

How do I do that

Vocaloid (vocaloid):

first, look at the left side of the equation 2(x-3) distribute the 2 and tell me what you get

OpenStudy (anonymous):

2x

Vocaloid (vocaloid):

not quite, try again hint: A(B-C) = A*B - A*C

OpenStudy (anonymous):

2x-6

Vocaloid (vocaloid):

right! so the left side of the equation is 2x - 6 and the right side of the equation is x + 5 2x - 6 = x + 5 can you try solving for x now?

OpenStudy (anonymous):

Do I subtract 6 to each side

OpenStudy (anonymous):

Or 5

Vocaloid (vocaloid):

add 6 to both sides, since we have -6 on the left side and we want to get rid of it

OpenStudy (anonymous):

2x=x+11 right?

Vocaloid (vocaloid):

right, now subtract x from each side

OpenStudy (anonymous):

X=11?

Vocaloid (vocaloid):

right, x = 11, only one more step to go! let's use the equation from before, BD=x+5 if x = 11, what is x + 5?

OpenStudy (anonymous):

16?

Vocaloid (vocaloid):

great! that's our answer

OpenStudy (anonymous):

Thank you! Could you help me with a few more?

Vocaloid (vocaloid):

sure

OpenStudy (anonymous):

how do I upload a pic on hete

Vocaloid (vocaloid):

"attach file"

OpenStudy (anonymous):

Where is that

Vocaloid (vocaloid):

right under the box where you type stuff in

OpenStudy (anonymous):

I'm on mobile I do not see it

Vocaloid (vocaloid):

Ah, I'm not familiar with the mobile site @pooja195 perhaps you can help?

OpenStudy (anonymous):

Is that person online

pooja195 (pooja195):

i dont use the mobile site o-o @paki @sleepyhead314

OpenStudy (anonymous):

Are they online

OpenStudy (sleepyhead314):

just a sec ^_^

jigglypuff314 (jigglypuff314):

Here, can you see this? There is a camera icon and paper clip button for you to attach pictures with :)

OpenStudy (anonymous):

I don't have that?

OpenStudy (anonymous):

??

OpenStudy (sleepyhead314):

uh oh which iphone are you using? (or itouch?) or have you updated the app recently?

OpenStudy (anonymous):

Do I have to be using the app I'm on the internet but I can get the app I'm on the IPhone 5s

OpenStudy (sleepyhead314):

ohhh yes! using the App would allow you to attach files :)

OpenStudy (anonymous):

I'm getting it

OpenStudy (anonymous):

I'm on the app now can you help me with a problem

OpenStudy (sleepyhead314):

sure ^_^

OpenStudy (anonymous):

@Vocaloid

OpenStudy (anonymous):

Ok here's the pic

OpenStudy (sleepyhead314):

I suggest attaching the pic with the app then going back to the in browser (because the app is still a little slow)

OpenStudy (anonymous):

Find the value of Y. Angle 1=3y+6

OpenStudy (sleepyhead314):

interior angle of a square is 90 degrees 3y + 6 = 90

OpenStudy (anonymous):

Is the answer 90?

OpenStudy (sleepyhead314):

not quite :) we have to solve for "Y" when given 3y + 6 = 90 you would first subtract 6 from both sides

OpenStudy (anonymous):

Or is it 28

OpenStudy (sleepyhead314):

that's correct :)

OpenStudy (anonymous):

Could you help me with a couple more

OpenStudy (sleepyhead314):

of course

OpenStudy (anonymous):

Find the measure of Angle 2

OpenStudy (sleepyhead314):

assuming this is a parallelogram?

OpenStudy (anonymous):

Yes

OpenStudy (sleepyhead314):

just a parallelogram or square rhomus kite? does it say anything else? :3

OpenStudy (anonymous):

It doesn't say but the topic is Special Parallelograms

OpenStudy (sleepyhead314):

ohhhh then it should definitely be a ... rhombus or something and the diagonals of a rhombus always bisect at 90 degree angles so if it really is a rhombus, then it should be 90 degrees ^_^

OpenStudy (anonymous):

The answer is 90?

OpenStudy (sleepyhead314):

I believe so, if there is no other information given

OpenStudy (anonymous):

It was right

OpenStudy (anonymous):

Find the measure of angle 2

OpenStudy (anonymous):

Forget the pic I just sent can you help with this one? RT=2x+4 SW=6x-20

OpenStudy (sleepyhead314):

diagonals in a rectangle are equal to each other so 2x + 4 = 6x - 20

OpenStudy (anonymous):

Add 6??

OpenStudy (sleepyhead314):

subtract 2x from both sides

OpenStudy (anonymous):

So 4=4x-20?

OpenStudy (sleepyhead314):

that's right! :) great job now add 20 to both sides

OpenStudy (anonymous):

24=4x

OpenStudy (sleepyhead314):

then divide both sides by 4

OpenStudy (anonymous):

X=6 ??

OpenStudy (sleepyhead314):

mhmm

OpenStudy (anonymous):

Find the measure of angle 1

OpenStudy (sleepyhead314):

assuming angle 1 equals angle 2 there are 360 degrees in a quadrilateral 360 = 2*104 + 2*(2[angle 1]) 360 = 208 + 4 [angle 1] now subtract 208 from both sides

OpenStudy (anonymous):

152=4?

OpenStudy (anonymous):

Divide by 4

OpenStudy (anonymous):

38??

OpenStudy (sleepyhead314):

that's what I got :)

OpenStudy (anonymous):

So it's right

OpenStudy (anonymous):

RZ=30 SW=5x-20

OpenStudy (sleepyhead314):

2(RZ) = (SW) 2(30) = 5x - 20

OpenStudy (anonymous):

Is the answer 16

OpenStudy (sleepyhead314):

yup

OpenStudy (anonymous):

It was wrong

OpenStudy (anonymous):

Find the measure of angle 4

OpenStudy (sleepyhead314):

it was wrong and yet you still ask me? xD perhaps you should try opening a new question to get someone to help you I'm not actually good at these :-/

OpenStudy (anonymous):

You've been getting them all right though

OpenStudy (sleepyhead314):

what I'm saying is idk how to do this one, sorry o-o